Two cars travel towards each other from points 400 km apart. The two cars meet in 4 hours. What is the average speed of each car if one car travels 20 km per hour faster than the other?
Answer by ikleyn(52781)

.

Let the slower car rate be x km/h;

then the faster car rate is  (x+20) km/h.


The total distance equation is

    4x + 4*(x+20) = 400  kilometers.


Simplify and find x

    4x + 4x + 80 = 400

    8x           = 400 - 80 fa

    8x           = 320

     x           = 320/8 = 40.


ANSWER.  The slower car rate is  40 km/h;  the faster car rate is 40+20 = 60 km/h.

Solved.

---------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.